1mole of oxygen initially at 7°c adiabatically compressed so that its pressure becomes 10 fold calculate its temperature after the compression, work done by the gas
1
Expert's answer
2019-04-03T09:14:40-0400

For adiabatic process:


"T_f=T_i(\frac{p_i}{p_f})^{\frac{1-\gamma}{\gamma}}"


"\frac{1-\gamma}{\gamma}=\frac{1-1.4}{1.4}=-\frac{2}{7}"


"T_i=273+7=280 K"

"\frac{p_i}{p_f}=\frac{1}{10}=0.1"

"T_f=280(0.1)^{-\frac{2}{7}}=540K"

The work done by the gas:


"W=\nu c_V (T_f-T_i)"

"W=(1)(21.1) (540-280)=5500 J=5.5 kJ"